how do I find the amount of interest earned by the following deposits ? $23000 at 12% compounded monthly for 21/2 years?
Answer by stanbon(75887)

$23000 at 12% compounded monthly for 21/2 years?
---
A(t) = P(1+(r/n))^(nt)
---
A(21/2) = 23000(1+(0.12/12))^(12*(21/2))
----
A(21/2) = 2300(1+0.1)^(126)
----
A(21/2) = 2300*164239.77
A(21/2) = 377,751,472.50
================
Cheers,
Stan H.
================